OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? A. 9,40, 42 B. 5,10,15 C. 6,8,10 D. 13,35,37 Please help (: will give a medal

OpenStudy (welshfella):

a right angled triangle will satisfy pythagoras theorem c^2 = a^2 + b^2 where c is the longest side

OpenStudy (welshfella):

try the first one if 42^2 = 40^2 + 9^2 then this is right angled triangle

OpenStudy (welshfella):

try each one till you get a hit